Two cars leave towns 510 kilometers apart at the same time and travel toward each other. One car's rate is 18 kilometers per hour more than the other's, If they meet in 3 hours, what is the rate of the faster car.
Answer by mananth(16949)

let speed of 1 car be x
other's speed will be x+18
They are travelling yowrads each other.
effective speed will be x+x+18
=2x+18
Distance = 510 km
speed = 2x+18
time = 3 hours
speed = distance / time
2x+18 = 510/3
2x+18= 170
2x= 170-18
2x= 152
x= 152 /2
x= 76 km/h
add 18 to x to find the speed of the other car